GP E = − ∑_{i≠j} \frac{Gm_{i}m_{j}}{r_{ij}}

KE = ∑_{i} \frac{1}{2}m_{i}v_{i}^{2}

Initial Energy = GP E + KE Final Energy = GP E + KE

\frac{ΔE}{E} = \frac{(Final Energy) − (Initial Energy)}{(Initial Energy)}